OpenStudy (anonymous):

How do you square root 13?

OpenStudy (dumbcow):

use a calculator.... the way it computes it is 13 is between 9 and 16 so sqrt13 is between 3 and 4 3.5^2 = 12.25 so sqrt13 is between 3.5 and 4 3.75^2 = 14.06 so sqrt13 is between 3.5 and 3.75 3.625^2 = 13.14 so sqrt13 is between 3.5 and 3.625 and so on until you get really close to the exact value

OpenStudy (anonymous):

Its 3.6 right? I had a feeling @dumbcow

OpenStudy (dumbcow):

sqrt13 = 3.60555

OpenStudy (anonymous):

Thanks (: @dumbcow

OpenStudy (dumbcow):

:)